If the birth rate of a population is b(t)= − 2200e


0.024t people per year and the death rate


is d(t)= − 1460e


0.018t people per year, find the area between these curves for 0 < t < 10.


What does this area represent? Also plot the curves using MATLAB.


1
Expert's answer
2021-12-14T12:55:13-0500

The death curve is blue and birth curve is red.



From graph, we can see that b(t) is greater than d(t)

Therefore

 "\begin{gathered}

\text { Area between the two curves }=\int_{0}^{10} b(t)-d(t) d t \\

=\int_{0}^{10} 2200 e^{0.024 t}-1460 e^{0.018 t} d t \\

=\left[\frac{2200 e^{0.024 t}}{0.024}-\frac{1460 e^{0.018 t}}{0.018}\right]_{0}^{10} \\

=\left[\frac{2200 e^{0.24}}{0.024}-\frac{1460 e^{0.18}}{0.018}\right]-\left[\frac{2200}{0.024}-\frac{1460}{0.018}\right] \approx 8868

\end{gathered}"

Since birth curve is greater than death curve we conclude that area between the two curves represents the increase in population between t=0 and t=10.